Two samples of the same radioactive nuclide are prepared. Sample G hadms twice the initial activity of sample H. How does the half-life of G compare with the half-life of H?
(A) it is two times larger than that of G
(B) it is thev same as that of G
(C) it is half of that of G
(D) it is larger than that of G 4 times.
1
Expert's answer
2021-02-17T11:10:06-0500

The relationship between the activity "R" and the number of nuclei in the sample, "N", can be written as follows:


"R=N\lambda,"

here, "\lambda" is decay constant.

Since, both samples have the same size (and, therefore, have the same initial number of nuclei), we can write:


"R_G=N_o\lambda_G,""R_H=N_o\lambda_H."

Dividing "R_G" by "R_H", we get:


"\dfrac{R_G}{R_H}=\dfrac{\lambda_G}{\lambda_H},""\dfrac{2R_H}{R_H}=\dfrac{\lambda_G}{\lambda_H},""\lambda_G=2\lambda_H."

The relationship between the half-life of a sample, "T_{1\/2}", and its decay constant "\lambda" can be written as follows:


"T_{1/2}=\dfrac{ln2}{\lambda}."

Then, we can write the half-life of each samples as follows:


"T_{1/2,G}=\dfrac{ln2}{\lambda_G},""T_{1/2,H}=\dfrac{ln2}{\lambda_H}."

Dividing "T_{1\/2,G}" by "T_{1\/2,H}" and substituting "\lambda_G" we get:


"\dfrac{T_{1/2,G}}{T_{1/2,H}}=\dfrac{\lambda_H}{\lambda_G}=\dfrac{\lambda_H}{2\lambda_H}=\dfrac{1}{2},""T_{1\/2,H}=2T_{1\/2,G}."

As we can see from calculations, the half-life of sample H is two times larger than that of G. Therefore, the correct answer is (A).

Answer:

(A) it is two times larger than that of G.